# VISTA: Knowledge-Driven Interpretable Vessel Trajectory Imputation via Large Language Models

### v1

@@ 4.3.2. Method Selector
This component operates similarly to the Behavior Estimator but omits contextual information during selection, as the selected behavior $v_{b}^{*}$ already encapsulates the relevant kinematic and intent information. Given the selected behavior $v_{b}^{*}$, it retrieves candidate functions through the behavior–function edges and scores them using the same prior construction as in Equations 9 and 10, yielding a ranked shortlist $\mathcal{C}_{f}^{K}$. It then builds the induced subgraph $\mathcal{G}_{\iota,f}^{K}=\mathcal{G}_{d}[\{v_{b}^{*}\}\cup\mathcal{C}_{f}^{K}]$, serialize it in DOT, and select the best method $v_{f}^{*}\in\mathcal{C}_{f}^{K}$ by leveraging the LLM’s analysis of graph evidence and compatibility with the chosen behavior $v_{b}^{*}$. Next, it executes $f\in v_{b}^{*}$ to obtain the reconstruction $\widehat{\mathbf{S}}_{\iota}^{k}$. Alongside the selection, it records graph support by citing the informative edges within $\mathcal{G}_{\iota,f}^{K}$, which forms the rationale $\mathcal{J}_{\iota}^{k,f}$.
@@ 4.3.3. Explanation Composer
Explanation composer delivers a human-friendly explanation designed for downstream tasks to internalize two kinds of domain knowledge: a Regulatory-rule cue (under which rules), and an Operational-protocol rationale (why it typically happens here). It analyzes a compact evidence view composed of the selected behavior $v_{b}^{*}$, the selected method $v_{f}^{*}$, static attributes $\mathcal{V}_{s}(\iota)$, spatial context $\sigma$, boundary patterns $(v_{b}^{-},v_{b}^{+})$, and the induced subgraph $\mathcal{G}_{d}[\mathcal{V}_{s}(\iota)\cup\{v_{b}^{*}\}\cup\{v_{f}^{*}\}]$. Specifically, the explanation consists of the following components:
@@ 4.4. Workflow Management Layer
While the previous sections establish how VISTA extracts knowledge and performs trajectory imputation, both stages rely heavily on LLM inference—an operation that is inherently slow and occasionally unreliable due to stochastic failures (e.g., malformed outputs, null responses, or execution errors) (Ahn and Yin, 2025; Tan et al., 2025; Agrawal et al., 2024). To ensure scalability and robustness when processing large AIS data, we design a workflow management layer that automates task scheduling, anomaly handling, and de-redundancy. This layer includes two complementary components: the SD-KG Construction Workflow Manager and the Trajectory Imputation Workflow Manager.
@@ 4.4.1. SD-KG Construction Workflow Manager
The SD-KG Construction Workflow Manager coordinates a two–stage parallel workflow, comprising extraction and de-redundancy. It consists: Stack-Based Scheduler, Anomaly Guard, and Deredundancy Processor.
Stack-Based Scheduler. All minimal segments are ordered by timestamp and pushed as job tuples onto a compute stack $\mathcal{S}_{c}$, while a second stack $\mathcal{S}_{d}$ buffers validated micro-batches before de-redundancy. At runtime, the scheduler pops up to $b$ jobs from $\mathcal{S}_{c}$ to form a micro-batch $\mathcal{B}_{t}$ and dispatches them to $|\mathcal{B}_{t}|$ worker threads. Each worker executes the extraction pipeline to produce a per-segment knowledge unit $u=(v_{s},v_{b},v_{f})$, incrementing the retry counter $c$ on failure and discarding the job when $c>\varrho_{c}$. Validated outputs are packed as a micro-batch $\widehat{\mathcal{B}}_{t}$ and pushed onto the second stack $\mathcal{S}_{d}$, which acts as a decoupling buffer between extraction and de-redundancy. This design preserves temporal order and metadata while isolating the downstream consolidation stage from upstream variability.
@@ 4.4.2. Trajectory Imputation Workflow Manager
Stack-Based Scheduler. All incomplete minimal segments are timestamp-sorted and pushed onto a stack $\mathcal{S}_{i}$ in descending order. The scheduler repeatedly pops a batch of size $b$ and executes the imputation pipeline in parallel. Each job runs the three modules in Section 4.3 to reconstruct the missing segment using the SD-KG. A failure counter is maintained per job; when it exceeds the retry threshold $\varrho_{i}$, the job is aborted and logged for offline diagnosis.
Anomaly Guard. Imputation anomalies primarily arise from response timeouts or invalid model outputs. Each job therefore undergoes two basic checks: 1) the response is non-empty (i.e., no request timeout), and 2) the selected behavior pattern $v_{b}^{*}$ and imputation function $v_{f}^{*}$ are correctly retrieved from the SD-KG and are executable. Jobs failing either check are re-queued for retry until the per-item limit $\varrho_{r}$ is reached, after which they are quarantined for offline inspection.
Finally, the selected imputation methods are executed to reconstruct the missing trajectory segments. The resulting trajectories, along with their explanations, are recorded for interpretability and traceability. Further details (e.g., prompt about redundancy and overall procedures) are provided in Appendix A.3
@@
Qwen-plus Qwen-plus Qwen-plus 1.817e-3 4.324e-2 1.123e-3 4.027e-2 0.2418 4.388e-3 1.455e-2 5.691e-3 2.052e-2 0.7945
Qwen-plus Qwen-plus Qwen-flash 5.706e-3 8.177e-2 5.578e-1 2.055e+1 2.8990 6.425e-3 2.957e-2 7.015e-1 1.311e+1 25.0557
Qwen-plus Qwen-flash Qwen-plus 2.872e-3 3.248e-2 5.818e-3 6.126e-2 0.5255 6.145e-3 2.371e-2 1.203e-2 6.505e-2 6.3462
Qwen-flash Qwen-plus Qwen-plus 1.902e-3 5.826e-2 1.216e-3 4.718e-2 0.2672 5.422e-4 1.7793e-2 8.960e-3 3.598e-2 1.8324
@@
Batch Size AISDK AISUS
@@ 5.1.2. Evaluation Metrics
For longitude and latitude, we report axis-wise Mean Absolute Error (MAE), a robust measure of average error, and Root Mean Squared Error (RMSE), which emphasizes large deviations, both expressed in degrees. For the joint spatial error, we use the Haversine distance (Wikipedia contributors, 2025b) with Earth’s mean radius R=6371R{=}6371 km and report the Mean Haversine Distance (MHD) in kilometers.
@@ 5.1.3. Baseline Methods
We compare VISTA with rule-based, deep learning-based, and LLM-based baselines. For rule based methods we include: Lin-ITP (Widyantara et al., 2023) linearly interpolates between anchors, Akima spline (Zaman et al., 2023) yields visually smooth, $C^{1}$-continuous paths that capture gentle curvature, and Kalman Filter (Wang et al., 2023) assumes a linear Gaussian state space with constant velocity and suits near linear motion. For deep based methods we include: MH-GIN (Liu et al., 2025) models multi scale features and cross attribute dependencies on a heterogeneous graph to impute all attributes in AIS data, and Multi-task AIS (Nguyen et al., 2018) is a recurrent framework with latent variables and AIS embeddings for noisy, irregular sampling. For LLM based methods we include: KAMEL (Musleh and Mokbel, 2023) treats completion as masked infilling with spatially aware tokenization; we also assess out of the box general models by using two families, Qwen (Qwen Team, 2025) and GLM 4.5 (Zeng et al., 2025), each with a lightweight and a full capacity variant with thinking mode enabled (Alibaba Cloud Team, 2025), yielding two pairs: Qwen-plus-th vs Qwen-flash-th and GLM-4.5-air-th vs GLM-4.5-th.
@@ 5.1.4. Experimental Settings
All experiments are conducted on a server with Intel Xeon Processor (Icelake) CPUs, 100GB RAM, and two NVIDIA A10 GPUs (each with 23GB memory). For data processing, each dataset is partitioned into 80% training and 20% testing. For comparability, all results reported below are computed on the held-out test split. To emulate realistic block missingness, trajectories are divided into minimal segments of length $m=20$, and each minimal segment is removed with probability $0.2$ (see Section 3.3). The mask configuration are shared for all methods. The more detailed experimental setting is provided in Appendix B.
Figure 7. Case study of VISTA.
@@ 5.2. Performance Evaluation
As shown in Table 2, which reports the results on AIS-DK and AIS-US across MAE, RMSE, and MHD, boldface denotes the best performance, while underlining indicates the second-best. Building on these results, we summarize four observations. 1) VISTA consistently outperforms all baselines, improving over the strongest baseline, MH-GIN, by 5%–94%. These results demonstrate its effective mining of AIS domain knowledge and productive use of underlying knowledge. 2) All models perform better on AIS-DK, and the inter-model spread is smaller than on AIS-US, consistent with shorter trajectory durations and smaller inter-sample time intervals in AIS-DK, which produce shorter missing gaps and an easier imputation setting. 3) Within the LLM group, the trajectory-oriented KAMEL performs markedly better than out-of-the-box general thinking-mode LLMs (Qwen-plus-th, Qwen-flash-th, GLM-4.5-th, GLM-4.5-air-th), underscoring the value of task-specific spatiotemporal tokenization and a “physical grammar” for trajectory imputation. Nevertheless, KAMEL still trails VISTA by a wide margin, highlighting the advantage of explicit, executable knowledge. The general models also show that “thinking” alone is insufficient, since capability strongly affects accuracy: higher-capacity variants outperform their lighter counterparts (Qwen-plus-th and GLM-4.5-th outperform Qwen-flash-th and GLM-4.5-air-th), and in our setting the GLM family tends to outperform Qwen (GLM-4.5-th and GLM-4.5-air-th outperform Qwen-plus-th and Qwen-flash-th), although all remain well below VISTA. 4) Deep-learning methods (Multi-task AIS, MH-GIN) generally outperform other kinds of baselines across metrics, reflecting the advantage of learning from AIS data; among them, MH-GIN is the strongest non-VISTA competitors overall, yet still trail VISTA by a wide margin.
@@ 5.3. Ablation Study
From Table 4, the Decision role is the most capacity-sensitive: downgrading it from Qwen-plus to Qwen-flash sharply increases MAE, RMSE, and MHD, showing that accurate behavior estimation and method selection demand stronger reasoning and contextual inference. In contrast, Analysis and Programming degrade more moderately under lighter models, as their semantic abstraction and method synthesis processes can still leverage the structured priors embedded in the SD-KG, which partially compensate for reduced model capacity. From Table 5, latency scales almost linearly with batch size, confirming that the workflow layer achieves efficient and contention-free parallel scheduling. Disabling de-redundancy slightly shortens runtime but significantly reduces size (e.g., $1,400\rightarrow 1,105$ on AIS-DK and $4,109\rightarrow 2,057$ on AIS-US), indicating a trade-off between speed and compact SD-KG.
5.4. Case Study

### v2

@@ 4.3.2. Method Selector
This component operates similarly to the Behavior Estimator but omits contextual information during selection, as the selected behavior $v_{b}^{*}$ already encapsulates the relevant kinematic and intent information. Given the selected behavior $v_{b}^{*}$, it retrieves candidate functions through the behavior–function edges and scores them using the same prior construction as in Equations 8 and 9, yielding a ranked shortlist $\mathcal{C}_{f}^{K}$. It then builds the induced subgraph $\mathcal{G}_{\iota,f}^{K}=\mathcal{G}_{d}[\{v_{b}^{*}\}\cup\mathcal{C}_{f}^{K}]$, serializes it in DOT, and selects the best method $v_{f}^{*}\in\mathcal{C}_{f}^{K}$ by leveraging the LLM’s analysis of graph evidence and compatibility with the chosen behavior $v_{b}^{*}$. Next, it executes $f\in v_{f}^{*}$ to obtain the reconstruction $\widehat{\mathbf{S}}_{\iota}^{k}$. Alongside the selection, it records graph support by citing the informative edges within $\mathcal{G}_{\iota,f}^{K}$, which forms the rationale $\mathcal{J}_{\iota}^{k,f}$.
@@ 4.3.3. Explanation Composer
The Explanation Composer delivers a human-friendly explanation that helps downstream tasks internalize two kinds of domain knowledge: a Regulatory-rule cue (under which rules), and an Operational-protocol rationale (why it typically happens here). It analyzes a compact evidence view composed of the selected behavior $v_{b}^{*}$, the selected method $v_{f}^{*}$, static attributes $\mathcal{V}_{s}(\iota)$, spatial context $\sigma$, boundary patterns $(v_{b}^{-},v_{b}^{+})$, and the induced subgraph $\mathcal{G}_{d}[\mathcal{V}_{s}(\iota)\cup\{v_{b}^{*}\}\cup\{v_{f}^{*}\}]$. Specifically, the explanation consists of the following components:
@@ 4.4. Workflow Management Layer
While the previous sections establish how VISTA extracts knowledge and performs trajectory imputation, both stages rely heavily on LLM inference—an operation that is inherently slow and occasionally unreliable due to stochastic failures (e.g., malformed outputs, null responses, or execution errors) (Ahn and Yin, 2025; Tan et al., 2025; Agrawal et al., 2024a). To ensure scalability and robustness when processing large AIS data, we design a workflow management layer that automates task scheduling, anomaly handling, and de-redundancy. This layer includes two complementary components: the SD-KG Construction Workflow Manager and the Trajectory Imputation Workflow Manager.
@@ 4.4.1. SD-KG Construction Workflow Manager
The SD-KG Construction Workflow Manager coordinates a two-stage parallel workflow—extraction and de-redundancy—through three modules: Stack-Based Scheduler, Anomaly Guard, and De-redundancy Processor.
Stack-Based Scheduler. All minimal segments are ordered by timestamp and pushed as job tuples onto a compute stack $\mathcal{S}_{c}$, while a second stack $\mathcal{S}_{d}$ buffers validated micro-batches before de-redundancy. At runtime, the scheduler pops up to $b$ jobs from $\mathcal{S}_{c}$ to form a micro-batch $\mathcal{B}_{t}$ and dispatches them to $|\mathcal{B}_{t}|$ worker threads. Each worker executes the extraction pipeline to produce a per-segment knowledge unit $u=(v_{s},v_{b},v_{f})$, incrementing the retry counter $c$ on failure and discarding the job when $c>\varrho_{c}$. Validated outputs are packed as a micro-batch $\widehat{\mathcal{B}}_{t}$ and pushed onto the second stack $\mathcal{S}_{d}$, which acts as a decoupling buffer between extraction and de-redundancy. This preserves temporal order while isolating consolidation from extraction-rate fluctuations.
@@ 4.4.2. Trajectory Imputation Workflow Manager
Stack-Based Scheduler. All incomplete minimal segments are timestamp-sorted and pushed onto a stack $\mathcal{S}_{i}$ in descending order. The scheduler repeatedly pops a batch of size $b$ and executes the imputation pipeline in parallel. Each job runs the three components in Section 4.3 to reconstruct the missing segment using the SD-KG. A failure counter is maintained per job; when it exceeds the retry threshold $\varrho_{i}$, the job is aborted and logged for offline diagnosis.
Anomaly Guard. Imputation anomalies primarily arise from response timeouts or invalid model outputs. Each job therefore undergoes two basic checks: 1) the response is non-empty (i.e., no request timeout), and 2) the selected behavior pattern $v_{b}^{*}$ is correctly retrieved from the SD-KG and the imputation function $v_{f}^{*}$ is executable. Jobs failing either check are retried up to $\varrho_{r}$ times before being quarantined for offline inspection.
Finally, the selected imputation methods are executed to reconstruct the missing trajectory segments. The resulting trajectories, along with their explanations, are recorded for interpretability and traceability. Further details (e.g., prompt about redundancy and overall procedures) are provided in Appendix A.3.
@@
Qwen-plus Qwen-plus Qwen-plus 1.817e-3 4.324e-2 1.123e-3 4.027e-2 0.2418 7.587e-3 2.050e-2 1.027e-2 2.942e-2 1.4532
Qwen-plus Qwen-plus Qwen-flash 5.706e-3 8.177e-2 5.578e-1 2.055e+1 2.8990 2.352e-2 3.418e-1 5.104e+0 1.501e+1 17.4384
Qwen-plus Qwen-flash Qwen-plus 2.872e-3 3.248e-2 5.818e-3 6.126e-2 0.5255 1.198e-2 4.637e-2 5.340e-2 8.913e-2 3.1527
Qwen-flash Qwen-plus Qwen-plus 1.902e-3 5.826e-2 1.216e-3 4.718e-2 0.2672 8.135e-3 2.263e-2 1.106e-2 3.354e-2 1.5847
@@
Batch Size AIS-DK AIS-US
@@ 5.1.2. Evaluation Metrics
For longitude and latitude, we report axis-wise Mean Absolute Error (MAE), a robust measure of average error, and Root Mean Squared Error (RMSE), which emphasizes large deviations, both in degrees. For the joint spatial error, we use the Haversine distance (Wikipedia contributors, 2025b) with Earth’s mean radius R=6371R{=}6371 km and report the Mean Haversine Distance (MHD) in kilometers.
@@ 5.1.3. Baseline Methods
We compare VISTA with rule-based, deep learning-based, and LLM-based baselines. For rule-based methods, we include: Lin-ITP (Widyantara et al., 2023) linearly interpolates between anchors, Akima Spline (Zaman et al., 2023) yields visually smooth, $C^{1}$-continuous paths that capture gentle curvature, and Kalman Filter (Wang et al., 2023) assumes a linear Gaussian state space with constant velocity and suits near-linear motion. For deep learning-based methods, we include: MH-GIN (Liu et al., 2025a) models multi-scale features and cross-attribute dependencies on a heterogeneous graph to impute all attributes in AIS data, and Multi-task AIS (Nguyen et al., 2018) is a recurrent framework with latent variables and AIS embeddings for noisy, irregular sampling. For LLM-based methods, we include: KAMEL (Musleh and Mokbel, 2023) treats completion as masked infilling with spatially aware tokenization; we also assess out-of-the-box general models by using two families, Qwen (Qwen Team, 2025) and GLM 4.5 (Zeng et al., 2025), each with a lightweight and a full-capacity variant with thinking mode enabled (Alibaba Cloud Team, 2025), yielding two pairs: Qwen-plus-th vs Qwen-flash-th and GLM-4.5-air-th vs GLM-4.5-th.
@@ 5.1.4. Experimental Settings
All experiments are conducted on a server with Intel Xeon Processor (Icelake) CPUs, 100 GB RAM, and two NVIDIA A10 GPUs (each with 23 GB memory). For data processing, each dataset is partitioned into 80% training and 20% testing. For comparability, all results reported below are computed on the held-out test split. To emulate realistic block missingness, trajectories are divided into minimal segments of length $m=20$, and each minimal segment is removed with probability $0.2$ (see Section 3.3). The mask configuration is shared across all methods. The more detailed experimental setting is provided in Appendix B.
@@ 5.2. Performance Evaluation
As shown in Table 1, which reports the results on AIS-DK and AIS-US across MAE, RMSE, and MHD, boldface denotes the best performance, while underlining indicates the second-best. Building on these results, we summarize four observations. 1) VISTA consistently outperforms all baselines, improving over the strongest baseline, MH-GIN, by 5%–91%. These results demonstrate that grounding LLM reasoning in data-verified SDK yields substantially more accurate imputation than any single-paradigm approach. 2) All models perform better on AIS-DK, and the inter-model spread is smaller than on AIS-US, consistent with shorter trajectory durations and smaller inter-sample time intervals in AIS-DK, which produce shorter missing gaps and an easier imputation setting. 3) Within the LLM group, the trajectory-oriented KAMEL performs markedly better than out-of-the-box general thinking-mode LLMs (Qwen-plus-th, Qwen-flash-th, GLM-4.5-th, GLM-4.5-air-th), underscoring the value of task-specific spatio-temporal tokenization and a “physical grammar” for trajectory imputation. Nevertheless, KAMEL still trails VISTA by a wide margin, highlighting the advantage of grounding LLM reasoning in a data-verified knowledge base (SD-KG). The general models also show that “thinking” alone is insufficient, since capability strongly affects accuracy: higher-capacity variants outperform their lighter counterparts (Qwen-plus-th and GLM-4.5-th outperform Qwen-flash-th and GLM-4.5-air-th), and in our setting, the GLM family tends to outperform its Qwen counterpart within each capacity tier (GLM-4.5-th vs. Qwen-plus-th, GLM-4.5-air-th vs. Qwen-flash-th), although all remain well below VISTA. 4) Among all baselines, MH-GIN achieves the best overall performance, followed closely by KAMEL and Multi-task AIS, reflecting the advantage of task-specific or data-driven modeling over rule-based and general-purpose LLM approaches. Nevertheless, MH-GIN still trails VISTA by a wide margin across all metrics. Beyond these quantitative results, a case study in Appendix C qualitatively illustrates how VISTA leverages SD-KG-driven reasoning to produce interpretable trajectory reconstructions with repair provenance.
@@
We also evaluate the reliability of the Imputation Method Builder’s code generation. As shown in Table 5, the first-pass success rate is 64.3% on AIS-DK and 29.5% on AIS-US. With the built-in verification-retry mechanism ($\varrho_{r}{=}3$), the success rate climbs steadily over successive attempts, reaching 77.4% and 70.2% after three attempts with an average of only 0.61 and 1.02 retries per segment, respectively. The lower first-pass rate on AIS-US reflects its longer trajectories and more complex traffic patterns near major U.S. coastal ports, which demand more sophisticated imputation functions; nevertheless, the retry mechanism narrows this gap substantially.
Table 5. Cumulative success rate (%) of LLM-generated imputation code over successive verification-retry attempts.
Dataset Attempt Avg. Retries
1st 2nd 3rd
AIS-DK 64.3 73.7 77.4 0.61
AIS-US 29.5 67.1 70.2 1.02
5.2.3. Scalability Analysis.
To evaluate how VISTA scales with data size, we run end-to-end experiments on subsets of 2,000, 5,000, and 10,000 trajectories for both AIS-DK and AIS-US. As shown in Table 6, the two datasets exhibit distinct scaling behaviors. On AIS-US, all five metrics remain nearly constant across the three scales (MHD varies within 1.41–1.45km), indicating that the SD-KG captures sufficient behavioral knowledge even from 2,000 trajectories; adding more data introduces neither improvement nor degradation, since the dense and repetitive port traffic near major U.S. coastal hubs provides adequate pattern coverage at smaller scales. On AIS-DK, performance is non-monotonic: MHD slightly increases from 0.34 at 2,000 to 0.40 at 5,000 before dropping to 0.24 at 10,000. This pattern suggests that at intermediate scales, the SD-KG absorbs more diverse but potentially conflicting behavior patterns from the heterogeneous Baltic and North Sea routes, temporarily increasing noise in candidate retrieval; as data grows further to 10,000, the statistical support for each pattern strengthens, enabling the knowledge base to disambiguate among candidates and produce more accurate imputation. Overall, VISTA maintains robust performance across all tested scales, confirming that the SD-KG construction and retrieval pipeline generalizes well without dataset-specific tuning.
Table 6. Scalability analysis with varying data sizes.
Dataset Trajec- tories MAE(Lat) RMSE(Lat) MAE(Lon) RMSE(Lon) MHD
AIS-DK 2,000 2.332e-3 5.838e-2 2.389e-3 4.988e-2 0.3435
5,000 2.850e-3 6.945e-2 2.247e-3 5.430e-2 0.3950
10,000 1.817e-3 4.324e-2 1.123e-3 4.027e-2 0.2418
AIS-US 2,000 7.504e-3 2.194e-2 1.035e-2 3.797e-2 1.4079
5,000 7.508e-3 1.968e-2 1.013e-2 2.972e-2 1.4310
10,000 7.587e-3 2.050e-2 1.027e-2 2.942e-2 1.4532
@@ 5.3. Ablation Study
[…]
5.4. Robustness Analysis
We evaluate VISTA’s robustness under two dimensions of data degradation: longer missing gaps (varying segment length $m$) and higher data loss (varying missing ratio). Both dimensions are external data conditions rather than internal design choices, and together they stress-test how well the SD-KG-grounded pipeline tolerates increasingly difficult imputation scenarios.
5.4.1. Varying Segment Length.
Table 7 reports performance under $m\in\{10,20,40\}$ with a fixed missing ratio of 20%. On AIS-DK, MHD increases from 0.06 at m=10m{=}10 to 0.24 at m=20m{=}20 and 0.59 at m=40m{=}40, remaining below 1km even when each missing gap spans 40 consecutive time steps. Notably, at m=10m{=}10 the MAE values drop to the $10^{-4}$-degree range (sub-100m accuracy), confirming that the SD-KG can almost perfectly reconstruct short gaps where boundary anchors tightly constrain the imputation. On AIS-US, performance is stable from m=10m{=}10 to m=20m{=}20 (MHD 1.40–1.45km) and increases moderately at m=40m{=}40 (MHD 2.46km, a 69% rise from m=20m{=}20), still remaining below 3km. The wider gap between AIS-DK and AIS-US at m=40m{=}40 is consistent with the dataset characteristics reported in Section 5: AIS-DK trajectories average 0.5 hours with relatively short Baltic and North Sea routes, so even 40-step gaps remain within the SD-KG’s pattern coverage; AIS-US trajectories average 2.8 hours with diverse maneuvers near major ports, giving the LLM a wider spatial extent to extrapolate over longer gaps.
Table 7. Imputation performance under varying segment lengths $m\in\{10,20,40\}$ with a fixed missing ratio of 20%.
Dataset $m$ MAE(Lat) RMSE(Lat) MAE(Lon) RMSE(Lon) MHD
AIS-DK 10 3.720e-4 5.264e-4 4.960e-4 7.980e-4 0.0571
20 1.817e-3 4.324e-2 1.123e-3 4.027e-2 0.2418
40 4.536e-3 8.908e-2 2.465e-3 4.683e-2 0.5949
AIS-US 10 7.188e-3 1.899e-2 9.969e-3 2.965e-2 1.3959
20 7.587e-3 2.050e-2 1.027e-2 2.942e-2 1.4532
40 1.221e-2 3.910e-2 1.887e-2 6.826e-2 2.4607
5.4.2. Varying Missing Ratio.
Table 8 reports performance under missing ratios of $\{10\%,20\%,30\%\}$ with a fixed segment length m=20m{=}20. On AIS-DK, VISTA is highly robust: MHD increases by only 17% from 0.22 at 10% to 0.25 at 30%, and MAE values remain nearly flat across all three ratios (e.g., MAE (Lat) varies within $1.3\text{e-}3$$2.1\text{e-}3$), indicating that per-segment imputation quality is largely unaffected by the proportion of missing data. On AIS-US, MHD rises steadily from 1.19 at 10% to 1.45 at 20% and 1.70 at 30%, a 42% total increase that remains moderate in absolute terms (<2<2km). Across both robustness experiments, VISTA maintains stable performance: AIS-DK remains within sub-kilometer accuracy under all tested conditions, while AIS-US shows gradual degradation that tracks the increasing difficulty but never exhibits catastrophic failure. The larger sensitivity on AIS-US reflects its longer trajectories and more complex traffic patterns, suggesting that enriching the SD-KG with more diverse behavioral patterns—e.g., through larger training sets or cross-region knowledge transfer—could further tighten the gap.
Table 8. Imputation performance under varying missing ratios $\{10\%,20\%,30\%\}$ with a fixed segment length m=20m{=}20.
Dataset Missing Ratio MAE(Lat) RMSE(Lat) MAE(Lon) RMSE(Lon) MHD
AIS-DK 10% 1.338e-3 3.653e-2 1.092e-3 3.114e-2 0.2167
20% 1.817e-3 4.324e-2 1.123e-3 4.027e-2 0.2418
30% 2.136e-3 7.138e-2 2.555e-3 1.298e-1 0.2529
AIS-US 10% 6.626e-3 2.001e-2 8.630e-3 2.953e-2 1.1932
20% 7.587e-3 2.050e-2 1.027e-2 2.942e-2 1.4532
30% 8.647e-3 2.555e-2 1.243e-2 3.913e-2 1.6977
To further illustrate how VISTA produces interpretable repair provenance for individual trajectory segments, we provide a detailed case study in Appendix C. We have also developed CLEAR (Liu et al., 2025b), an interactive demo platform built on VISTA that enables non-expert users to explore, complete, and annotate vessel trajectories through a knowledge-graph-driven interface.